OpenStudy (angelwings996):

Describe the vertical asymptote(s) and hole(s) for the graph of y = (x + 2)(x + 4) _______________ (x + 4)(x + 1)

OpenStudy (angelwings996):

a. asymptotes: x = -1 and hole: x = 4 b. asymptotes: x = 2 and hole: x = -1 c. asymptotes: x = -1 and hole: x = -4 d. asymptotes: x = 1 and hole: x = 4

OpenStudy (mathstudent55):

Notice that the (x + 4) in the numerator and the (x + 4) in the denominator cancel out. That means that this function has a graph just like y = (x + 2)/(x + 1). This has an asymptote at x = -1. What does the (x + 4) in the numerator and denominator do? It creates a hole at x = -4. Other than that, this is exactly like y = (x + 2)/(x + 1)

OpenStudy (angelwings996):

Okay thanks(: Do you think you could help me with one more ? @mathstudent55

OpenStudy (mathstudent55):

The 5 vs 1 in the numerator causes the shape of teh hyperbola to become flatter.